OpenStudy (anonymous):

If five cards are chosen at random from a standard deck of playing cards, how many different ways are there to draw the five cards if at least three cards are a jack, queen or a king?

OpenStudy (kropot72):

There are 4 each of Jacks, Queens and Kings making 12 cards. This leaves 52 - 12 = 40 other cards. If 3 of the 5 drawn cards are Jack, Queen or King the number of combinations is 12C3. Each of these combinations can be taken with each of the 40C2 combinations of the other cards, making a total of 12C3 * 40C2. If 4 of the 5 drawn cards are Jack, Queen or King the number of combinations is 12C4. Each of these combinations can be taken with each of the 40C1 combinations of the other cards, making a total of 12C4 * 40C1. If 5 of the 5 drawn cards are Jack, Queen or King the number of combinations is 12C5. Therefore the number of ways to draw 5 cards with at least 3 a Jack, Queen or King is: \[(12C3\times40C2)+(12C4\times40)+12C5=can you\ calculate?\]

OpenStudy (anonymous):

distribute?

OpenStudy (kropot72):

There are various ways to calculate. My preference is to calculate each product and then add the results.

OpenStudy (anonymous):

SO 12 times 40

OpenStudy (anonymous):

960?

OpenStudy (kropot72):

\[12C3\times40C2=\frac{12!\times40!}{3!\times9!\times2!\times38!}=\frac{12\times11\times10\times40\times39}{3\times2\times2}=you\ can\ calculate\]

OpenStudy (anonymous):

2,059,200/12

OpenStudy (kropot72):

Correct! Now do the division and you have the first product.

OpenStudy (anonymous):

which is 171,600

OpenStudy (anonymous):

thats how many ways?

OpenStudy (kropot72):

Correct again! Now for the next product: \[12C4\times40=\frac{12\times11\times10\times9\times40}{4\times3\times2}=you\ can\ calculate\]

OpenStudy (anonymous):

ohh so 171,600

OpenStudy (anonymous):

isnt the final answer?

OpenStudy (kropot72):

No it's not.

OpenStudy (anonymous):

19,800?

OpenStudy (kropot72):

Correct! Now we find the value of 12C5 and add this value to 171600 + 19800. \[12C5=\frac{12\times11\times10\times9\times8}{5\times4\times3\times2}=you\ can\ calculate\]

OpenStudy (anonymous):

95,040 over 120

OpenStudy (anonymous):

792??

OpenStudy (kropot72):

Correct! So, finally, we have to add: 171600 + 19800 + 792 = ? ways

OpenStudy (anonymous):

192,192

OpenStudy (kropot72):

Correct again. Good work!

OpenStudy (anonymous):

thanks

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

do you do tutoring

OpenStudy (kropot72):

I tutor only on OpenStudy.

OpenStudy (anonymous):

ok ive been struggling

OpenStudy (kropot72):

Well I'll try to help when I'm logged in :)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

the words work can be arranged in how many different ways

OpenStudy (anonymous):

24?

OpenStudy (anonymous):

same with the word KINNIKINNIK

OpenStudy (kropot72):

The letters in the word 'work' taken all at a time can be arranged in 4! = 24 ways. There are 11 letters in the word KINNIKINNIK, K is repeated 3 times, I is repeated 4 times and N is repeated 4 times. The number of different ways of arranging the word KINNIKINNIK is: \[\frac{11!}{3! \times4! \times4!}=\frac{11\times10\times9\times8\times7\times6\times5}{3\times2\times4\times3\times2}=you\ can\ calculate.\]

OpenStudy (anonymous):

1,663,200/144?

OpenStudy (anonymous):

= 11,550

OpenStudy (kropot72):

The reason for dividing by \[3! \times4! \times4!\] is to deal with the repeated letters which can be arranged a number of times without making a new arrangement of the total letters. Your result 11,550 ways is correct!

OpenStudy (anonymous):

thats the final answer?

OpenStudy (kropot72):

Yes, that's it!

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

does P (6,2) = 30?

OpenStudy (kropot72):

You're welcome. Please post any more questions as new ones, the reason being that I must log out now. BTW 30 is correct.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

appreciate it your like the best on here lol

OpenStudy (kropot72):

np